# Ripple: Accelerating LLM Inference on Smartphones with Correlation-Aware Neuron Management

### v1

@@ 4. Offline Correlation-Aware Clustering
Figure 8. An example of neuron placement search algorithm. Based on co-activation patterns, the correlation of all neuron pairs is ranked. These pairs are then sequentially retrieved, finally constructing a linking that incorporates all neurons.
@@ 4.2. Step 2: Graph-Based Problem Abstraction
Hamilton Pathfinding. The Hamiltonian path ensures that all neurons stored within the flash are involved, while minimizing the path length maximizes the likelihood of clustering co-activated neurons together. We prove that this problem can be further reduced to the classical Traveling Salesman Problem (TSP) (tsp, ) as follows:
Lemma 4.1.
The global shortest Hamiltonian path problem (i.e., finding the shortest path between any pair of nodes) can be reduced to the TSP in polynomial time.
Proof.
Define Problem A as: Find the shortest Hamiltonian path in a complete graph from node $x$ to node $y$, denoted as $f(x,y)$. By iterating over all possible node pairs $x$ and $y$, it is evident that the global shortest Hamiltonian path problem can be reduced to solving Problem A for all node pairs. This can be achieved in $O(n^{2})$ time using dynamic programming, where $n$ is the number of nodes in the graph.
On the other hand, the TSP is defined as: Find the shortest Hamiltonian cycle in a complete graph starting and ending at a node $x$, denoted as $g(x)$. The relationship between Problem A and the TSP can be established as follows:
(6) $g(x)=\min_{\begin{subarray}{c}i=1\\ i\neq x\end{subarray}}^{n}f(x,i)+\text{dist}(x,i)$
This expression shows that Problem A can be further reduced to the TSP in $O(n)$ time by minimizing over $n-1$ possible nodes. Therefore, the global shortest Hamiltonian path problem can be reduced to the TSP in polynomial time. ∎
@@ 4.3. Step 3: Heuristic-Driven Greedy Algorithm
Although the problem can be reduced into the classical TSP, which is known to be NP-hard (np-hard, ), finding an optimal solution in polynomial time is generally infeasible. We propose a heuristic algorithm that searches for the optimal neuron placement greedily, as shown in Figure 8
@@ 7. Related Works
Model Parameter Reduction. To alleviate the memory and computational burdens of LLM execution, substantial efforts have focused on the reduction of model parameters, with two principal approaches emerging. The first, Model Pruning (pruning-survey-1, ; pruning-survey-2, ), seeks to reduce the number of model parameters while ensuring minimal performance degradation. Several works (pruning-static-group, ; pruning-static-filter, ; pruning-static-sparse, ; pruning-static-element, ; pruning-static-deep-compression, ) have explored static pruning, where parameters are pruned offline, prior to inference. In contrast, dynamic sparsity methods (pruning-dynamic-rl, ; pruning-dynamic-sft, ; pruning-dynamic-rethink, ; pruning-dynamic-surgery, ; pruning-dynamic-low-rank, ; pruning-dynamic-movement, ) determines which parameters to prune during runtime, enabling seamless integration with training or inference. Different from these pruning techniques, Ripple exploits the Activation Sparsity inherent in LLMs, retaining all model parameters but enhancing resource efficiency by selectively activating only a subset. This approach preserves the model’s generalization ability, particularly critical in the context of LLMs.
@@
Activation Sparsity Application. Many recent works have begun leveraging activation sparsity to reduce the resource demands of LLM inference. For instance, Deja Vu (deja-vu, ) pioneered a predictor-based approach for sparsity-based LLM inference, greatly reducing inference latency. Building upon this, Powerinfer (powerinfer, ) exploits this property to enable LLM execution on consumer-grade GPUs by offloading model parameters to CPU. Particularly in mobile scenarios, LLM in a Flash (llm-flash, ) first proposes using flash on smartphones for model offloading. Powerinfer2 (powerinfer-2, ) extends this approach further, serving a 47B LLM on a smartphone. However, these methods primarily concentrate on optimizing DRAM management and overlapping computation with data transfers, achieving only limited bandwidth improvements. Ripple complements these efforts by directly enhancing the neuron transfer bandwidth, an optimization that can integrate with existing techniques to accelerate LLM inference on smartphones.

### v2

@@ 4. Offline Correlation-Aware Clustering
Figure 8. A search algorithm example. Based on co-activation patterns, neuron pairs are ranked by correlation and sequentially retrieved to form neuron linkings. During construction, the root of each linking (R) and the neighbor counts of each neuron (N) are recorded. By greedily merging the nearest linkings, a complete linking that encompasses all neurons is obtained.
@@ 4.2. Step 2: Graph-Based Problem Abstraction
Hamilton Pathfinding. The Hamiltonian path ensures that all neurons stored within the flash are involved, while minimizing the path length maximizes the likelihood of clustering co-activated neurons together. To find the optimal neuron placement, we have to search for the global shortest Hamiltonian path (i.e., finding the shortest path between any pair of nodes). We find this problem can be further reduced to the classical Traveling Salesman Problem (TSP) (tsp, ).
@@ 4.3. Step 3: Heuristic-Driven Greedy Algorithm
Although the problem can be reduced into the classical TSP, which is known to be NP-hard (np-hard, ), finding an optimal solution in polynomial time is generally infeasible. We propose a heuristic algorithm that searches for the optimal neuron placement greedily, as shown in Figure 8.
@@ 7. Related Works
Model Parameter Reduction. To alleviate the memory and computational burdens of LLM execution, substantial efforts have focused on the reduction of model parameters, with two principal approaches emerging. The first, Model Pruning (pruning-survey-1, ; pruning-survey-2, ), seeks to reduce the number of model parameters while ensuring minimal performance degradation. Several works (pruning-static-group, ; pruning-static-filter, ; pruning-static-sparse, ; pruning-static-element, ; pruning-static-deep-compression, ) have explored static pruning, where parameters are pruned offline, prior to inference. In contrast, dynamic sparsity methods (pruning-dynamic-rl, ; pruning-dynamic-sft, ; pruning-dynamic-rethink, ; pruning-dynamic-surgery, ; pruning-dynamic-low-rank, ; pruning-dynamic-movement, ) determines which parameters to prune during runtime, enabling seamless integration with training or inference. Different from these pruning techniques, Ripple exploits the Activation Sparsity inherent in LLMs, retaining all model parameters but enhancing resource efficiency by selectively activating only a subset. This approach preserves the model’s generalization ability, particularly critical for LLMs.
@@
Activation Sparsity Application. Many recent works have begun leveraging activation sparsity to reduce the resource demands of LLM inference. For instance, Deja Vu (deja-vu, ) pioneered a predictor-based approach for sparsity-based LLM inference, greatly reducing inference latency. Building upon this, Powerinfer (powerinfer, ) exploits this property to enable LLM execution on consumer-grade GPUs by offloading model parameters to CPU. Particularly in mobile scenarios, LLM in a Flash (llm-flash, ) first proposes using flash on smartphones for model offloading. Powerinfer2 (powerinfer-2, ) extends this approach further, serving a 47B LLM on a smartphone. However, these methods primarily concentrate on optimizing DRAM management and overlapping computation with data transfers, achieving only limited bandwidth improvements. Ripple complements these efforts by directly enhancing the neuron transfer bandwidth, an optimization that can integrate with existing techniques to fast LLM inference on smartphones.